dvired(1) teTeX dvired(1) NAME dvired−printdvi-files SYNOPSIS dvired[options]file DESCRIPTION dviredbehaveslikedvips,exceptforthefactthattwologicalpagesareputontoeachphysical sheetofpaper. Foroptions,seedvips(1).Thisprogramonlyinterpretsthefollowingoptions: −ofile−redirectoutputintofile. −Pprinter−directoutputtoprinter. −f−directoutputtostdout. Anyotheroptionispasseddirectlytodvips. EXAMPLES ItisassumedthatthePRINTER-variableisset. dvired−Plwfoo−sendoutputtoprinterlw. dvired−ofoo.psfoo−sendoutputtofilefoo.ps. dvired−pp4-7foo−send4pages(reducedonto2sheets)offoototheprinter. dviredfoo−f|ghostview−−previewoutputwithghostview. SEEALSO dvips(1) BUGS dviredisfixedonA4papersize.Ifyouwanttouseitwithotherpaper,youhavetochangethe settingswithinthesource. AUTHOR ThomasEsser(te@informatik.uni−hannover.de) teTeX 09/94 1